OpenStudy (anonymous):

How do you find the derivative of (t-2)^3 (t-6) ?

OpenStudy (jhannybean):

1 possible way is to expand \(\ (t-2)^3\), multiply it into (t-6) and then take the derivative of the expansion.

OpenStudy (anonymous):

just expand out the whole thing to make it easier x`(t)=4*t^3-30*t^2+72*t-54 x`(t)=(2*(2*t-3))*(t-3)^2

OpenStudy (jhannybean):

\[\ (t-2)^3 = t^3-6t^2 +12t - 8 \]\[\ (t-2)^3(t-6) = t^4-12t^3+48t^2-80t+48\]

OpenStudy (jhannybean):

And then just take the derivative of this \[\frac{d}{dt} (t^4-12t^3+48t^2-80t+48)\]

OpenStudy (jhannybean):

And you know how to take derivatives, right?

OpenStudy (anonymous):

yeah I'm just kind of new at this. I'm used to taking derivatives of let's say (t-2)^3 the extra (t-6) just throws me off. The chain rule is not needed right?

OpenStudy (jhannybean):

Nope, you can just take the derivative of the expansion as i showed above.

OpenStudy (jhannybean):

\[\ nt^{n-1}\]

OpenStudy (unklerhaukus):

\[y(t)=(t-2)^3 (t-6)\] The product rule \[y'(t)=\frac{\mathrm d}{\mathrm dx}\Big((t-2)^3\Big)(t-6)+(t-2)^3\frac{\mathrm d}{\mathrm dx}\Big(t-6\Big)\] Now use the cain rule to find \(\frac{\mathrm d}{\mathrm dx}\Big((t-2)^3\Big)\)

OpenStudy (unklerhaukus):

chain*

OpenStudy (jhannybean):

Oh that's another method! Nice.

OpenStudy (unklerhaukus):

The expansion method should give an equivalent result, but I think the product rule is a bit simpler for this question

OpenStudy (jhannybean):

Yeah, I agree.

OpenStudy (anonymous):

ok now I just have to stare at it to understand it. Thanks for your help :)

OpenStudy (unklerhaukus):

remember that the product rule is just \[y(x)=f(x)\cdot g(x)\\ y'(x)=f'g+fg'\] and the chain rule is\[y(x)=f(g(x))\\y'(x)=f'(g(x))\cdot g'\]

OpenStudy (mathstudent55):

Product rule & chain rule.

OpenStudy (mathstudent55):

\(y = uv\) \(y' = uv' + vu'\) \(y = (t - 2)^3(t - 6) \) \(y' = (t - 2)^3 (t - 6)' + (t - 6)[(t - 2)^3]'\) \(y' =(t - 2)^3 (1) + (t - 6)[3(t - 2)^2(1)]\) \(y' = (t - 2)^3 + 3(t -6)(t - 2)^2\) \(y' = (t - 2)^2 [(t - 2) + 3(t - 6)] \) \(y'= (t - 2)^2[t - 2 + 3t - 18]\) \(y' = (t - 2)^2 (4t -20)\) \(y' = (t - 2)^2 (4)(t - 5)\) \(y' = 4(t - 2)^2(t - 5) \)

OpenStudy (anonymous):

ok so is the derivative of (t-2)^3 this --> 3(t-2)^2 (1)

OpenStudy (mathstudent55):

Yes.

OpenStudy (jhannybean):

Yup

OpenStudy (anonymous):

ok I'm confused before I came here I asked this question on yahoo answers and someone said the derivative was 3(t-2)^2 (t-6)+ (t-2)^3 . Is it the same thing as the answer on here?

OpenStudy (jhannybean):

they half-used the product rule. It is incorrect.

OpenStudy (anonymous):

oh ok thank you so much for your help :D I finally understand it

OpenStudy (mathstudent55):

@endoftime14 Notice that the answer you got from yahoo is the same as the 6th line of my answer above (with the order of the terms switched). The answer is correct, but I just simplified it more.